# Inferring the Equation of State from Neutron Star Observables via Machine Learning

### v1

@@ 1 Introduction
This paper investigates the potential of machine learning, particularly the symbolic regression method, to uncover the hidden relationships between neutron star properties and the equation of state. For that purpose, we use a variety of neutron star equations of state, encompassing different compositions and several modelling frameworks, including meta-models. The EoS collection includes non-linear (NL) nucleonic and hyperonic (NL-hyp) models, models taken from the CompOSE database (CompOSE), models constructed from the extrapolation of the speed of sound (CSE), piecewise-polytropic (PWP) models, density-dependent models obtained from Bayesian inference (DDB), and models from a spectral representation (SR), providing a comprehensive foundation for our investigation. Exploring the relationships between various NS properties through correlation analysis offers a compelling approach to understand the dense matter EoS [53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63]. Specifically, in this work, we applied a symbolic regression method to a wide-ranging dataset to investigate: (i) the relation between the NS maximum mass and EoS properties at different energy densities, and (ii) the interplay between neutron star properties and the central energy density, and pressure for the neutron star masses in the range of 1-2.
@@ 2.2 Sampling for symbolic regression
A comprehensive outline of our symbolic regression methodology is presented in Fig. 1. To initiate the process, we extract the characteristic vectors indicated as and the corresponding target vector indicated as from our data set. Subsequently, we enter the next phase, where we randomly select a subset of characteristics from the original characteristics , as well as a single random target from the set of original targets. These chosen random features and targets are then fed into the PySR algorithm. For effective hyperparameter tuning of the PySR algorithm, including parameters like the number of iterations and operators, we employ two optimization strategies: (i) Bayesian optimization and (ii) Grid search. The optimal equation is selected using a dual approach involving both techniques. The equation with the highest Pearson correlation coefficient and smaller relative error (RE in %) between and is identified as the best equation. This whole process is iterated over 500,000 times.
@@ 3 Results and Discussion
As previously discussed in Section 2.1, we use a diverse set of EoSs for cold dense neutron star matter, encompassing different compositions and several models including meta-models. These collections of EoS correspond to NL, NL-hyp, CompOSE, CSE, PWP, DDB, and SR models. Before embracing on our results, it is essential to examine the diversity of our EoSs concerning the neutron star properties. In Figure 2, on the left, we display the 90% confidence interval (CI) for the radius of neutron stars across masses from 1 to the maximum mass obtained for each of the EoS set. This is compared with a range of astrophysical data. The shaded grey regions show the 90% (solid line ) and 50% (dashed line) confidence intervals of the LIGO / Vigo constraints, derived from the binary components of GW170817 [83]. Moreover, the 1 (68%) confidence intervals for the 2D posterior distribution in the mass-radius region for the millisecond pulsar PSR J0030+0451 are depicted in pastel blue and soft green [22, 21]. For the pulsar PSR J0740+6620, the confidence intervals, based on NICER X-ray data, are shown in blush pink. The most recent mass-radius data for the closest pulsar, PSR J0437-4715, is also included (orange) [84]. The figure demonstrates that all our EoS collections are in harmony with the constraints imposed by astrophysical observations. The CompOSE set covers the widest range in mass and radius as it contains EoS with different compositions, and models, relativistic and non-relativistic, and most of the EOS were constrained using properties of finite nuclei, only sensitive to densities close to saturation density. On the right, the figure shows the mass against the dimensionless tidal deformability for all EoS collections. The blue bar indicates the tidal deformability constraints for the NS with mass 1.36, as reported by Abbot et al.[18]. The figure clearly shows that the various EoS sets examined in this study encompass a wide range of NS properties, whether in terms of mass-radius or mass-tidal deformability, all while adhering to astrophysical observation constraints.
@@ 3.1.1 Maximum NS Mass
We aim to determine the relationship between the NS maximum mass and EoS properties such as energy density and pressure across different baryon densities in a model-independent manner using our diverse sets of EoS. In the left panel of Figure 3, we display the values of the Pearson’s correlation coefficient between pressure (P) and maximum mass of NS () as a function of energy density, for the nucleonic EoSs (purple line) that correspond to the NL and DDB datasets, and compare with those obtained for all the EoSs (blue dashed line). We observe that the correlation becomes stronger within a specific range of energy densities for nucleonic EoSs. This correlation is weak and negative at low energy densities and then increases up to () at 800 MeV fm-3 and tends to saturate. For the combined EoS results (All), the correlation peaks at 600 MeV·fm-3 with a Pearson coefficient of 0.85. In contrast, the individual sets exhibit correlations of [0.99, 0.97, 0.94, 0.88, 0.97, 0.64, 0.75] for the DDB, NL, NL-hyp, SR, CompOSE, CSE, and PWP models, respectively. Beyond 600 MeV·fm-3, the overall correlation declines, reaching 0.8 at 800 MeV·fm-3, while the individual correlations adjust to [0.99, 0.94, 0.87, 0.78, 0.95, 0.37, 0.66] for the same models. As might be expected, the agnostic descriptions show the smallest correlations, with PWP models still showing a correlation above 70%, 75% at MeV·fm-3. At MeV·fm-3, the correlation suffers an overall decrease which is particularly strong for the agnostic descriptions: for CSE it goes below 50%. Given these findings, we applied the symbolic regression method, as described in Sec. 2.2, to further investigate potential relationships between NS maximum mass with pressure and baryon density at a given energy density. This analysis helped us establish a relationship for mixed EoS models at an energy density of 800 MeV fm-3. The obtained equation is
@@
where represents the pressure and is the baryon density at an energy density of 800 MeV.fm-3. This correlation is shown in the right panel of Figure 3. Overall, it illustrates that over 90% of the models within the mixed set adhere closely to the derived equation, highlighting its robustness and practical applicability. The lower panel displays the fit error () as a function of , which remains below 4%, indicating the reliability of our fitted function. This relationship, being largely model-independent and reliable, will be employed for realistic applications in the following.
@@
We used the PyMultiNest sampler with 1500 live points on the Deucalion high performance computing (HPC) system at the University of Minho, Portugal, using 4 nodes with 128 CPUs each. The summary of CPU times for both analyses is shown in Table 1. It shows a dramatic reduction in sampling time for the scenario where the TOV equation was bypassed. This resulted in a reduction in CPU time of more than a factor of seven. We then examined the posterior distributions for the NS mass-radius relationships for both cases. The comparison involved around 8,000 EoSs for each case, as can be seen from Table 1.
@@ 3.1.2 Mass-Radius-Tidal deformability
The equations of state can be immediately obtained, from the observed mass-radius-tidal deformability, by using separately Eqs. (3) & (4)(SRM-MR) and Eqs. (5) & (6)(SRM-M). To test the reliability of these fits, we selected from each of the different sets shown in Figure 2, 100 MR and 100 M curves to reconstruct the EoSs using the relationships in Eqs. (3-6). These 700 reconstructed EoSs are compared to the original EoSs (All) in Figure 5. The 90% confidence interval for the variation of pressure with energy density reconstructed using SRM-MR (light pink) and SRM-M (cyan) are compared to the original EoSs (black). The reconstructed bands show a good overlap with the original one in the range MeV/fm3. It may be noted that the deviation for the reconstructed EoSs increases somewhat with density. We have also repeated the calculations for different sets of randomly selected EoSs, the results remained more-or-less the same.
@@ 4 Conclusions
We have obtained the expressions relating the equations of state to the NS observables such as the radius, and tidal deformability in the mass range of 1-2 . The EoSs derived using these expressions (Eqs. (3-6)) using NS observables are found to have good overlap with the actual EoSs.

### v2

@@ 1 Introduction
This paper investigates the potential of machine learning, particularly the symbolic regression method, to uncover the hidden relationships between neutron star properties and the equation of state. For that purpose, we use a variety of neutron star equations of state, encompassing different compositions and several modelling frameworks, including meta-models. The EoS collection includes non-linear (NL) nucleonic and hyperonic (NL-hyp) models, models taken from the CompOSE database (CompOSE), models constructed from the extrapolation of the speed of sound (CSE), piecewise-polytropic (PWP) models, density-dependent models obtained from Bayesian inference (DDB), and models from a spectral representation (SR), providing a comprehensive foundation for our investigation. Exploring the relationships between various NS properties through correlation analysis offers a compelling approach to understand the dense matter EoS [53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65]. Specifically, in this work, we applied a symbolic regression method to a wide-ranging dataset to investigate: (i) the relation between the NS maximum mass and EoS properties at different energy densities, and (ii) the interplay between neutron star properties and the central energy density, and pressure for the neutron star masses in the range of 1-2.
@@ 2.2 Sampling for symbolic regression
A comprehensive outline of our symbolic regression methodology is presented in the Figure 1. To initiate the process, we extract two types of vectors from our dataset: feature vectors and target vector . In these vectors, and represents the total number of feature and target quantity, respectively. We then enter the sampling phase, where we randomly select a subset of n feature () from the original characteristics (). Along with these, we also choose a single random target from the original set of targets . These randomly sampled combinations of features and target are fed into the PySR algorithm to discover mathematical relationships between them. For effective hyperparameter tuning of the PySR algorithm, including parameters like the number of iterations and operators, we employ two optimization strategies: (i) Bayesian optimization and (ii) Grid search. The optimal equation is selected using a dual approach involving both techniques. The equation with the highest Pearson correlation coefficient and smallest relative error (RE in %) between and is identified as the best equation. This entire process is iterated over 1/2 million times to explore different feature combinations and identify the most robust mathematical expressions that relate our EoS quantities to the NS properties. We performed this analysis in two directions: first, using EoS quantities as features and NS properties as targets, and then reversed the approach by using NS properties as features and EoS quantities as targets to explore bidirectional relationships in the data.
@@ 3 Results and Discussion
As previously discussed in Section 2.1, we use a diverse set of EoSs for cold dense neutron star matter, encompassing different compositions and several models including meta-models. These collections of EoS correspond to NL, NL-hyp, CompOSE, CSE, PWP, DDB, and SR models. Before embracing on our results, it is essential to examine the diversity of our EoSs concerning the neutron star properties. In Figure 2, on the left, we display the 90% confidence interval (CI) for the radius of neutron stars across masses from 1 to the maximum mass obtained for each of the EoS set. This is compared with a range of astrophysical data. The shaded grey regions show the 90% (solid line ) and 50% (dashed line) confidence intervals of the LIGO / Virgo constraints, derived from the binary components of GW170817 [85]. Moreover, the 1 (68%) confidence intervals for the 2D posterior distribution in the mass-radius region for the millisecond pulsar PSR J0030+0451 are depicted in pastel blue and soft green [22, 21]. For the pulsar PSR J0740+6620, the confidence intervals, based on NICER X-ray data, are shown in blush pink. The most recent mass-radius data for the closest pulsar, PSR J0437-4715, is also included (orange) [86]. The figure demonstrates that all our EoS collections are in harmony with the constraints imposed by astrophysical observations. The CompOSE set covers the widest range in mass and radius as it contains EoS with different compositions, and models, relativistic and non-relativistic, and most of the EOS were constrained using properties of finite nuclei, only sensitive to densities close to saturation density. On the right, the figure shows the mass against the dimensionless tidal deformability for all EoS collections. The blue bar indicates the tidal deformability constraints for the NS with mass 1.36, as reported by Abbot et al.[18]. The figure clearly shows that the various EoS sets examined in this study encompass a wide range of NS properties, whether in terms of mass-radius or mass-tidal deformability, all while adhering to astrophysical observation constraints.
@@ 3.1.1 Maximum NS Mass
We aim to determine the relationship between the NS maximum mass and EoS properties such as energy density and pressure across different baryon densities in a model-independent manner using our diverse sets of EoS. In the left panel of Figure 3, we display the values of the Pearson’s correlation coefficient between pressure and maximum mass of NS () as a function of energy density, for the nucleonic EoSs (purple line) that correspond to the NL and DDB datasets, and compare with those obtained for all the EoSs (blue dashed line). We observe that the correlation becomes stronger within a specific range of energy densities for nucleonic EoSs. This correlation is weak and negative at low energy densities and then increases up to () at 800 MeV fm-3 and tends to saturate. For the combined EoS results (All), the correlation peaks at 600 MeV·fm-3 with a Pearson coefficient of 0.85. In contrast, the individual sets exhibit correlations of [0.99, 0.97, 0.94, 0.88, 0.97, 0.64, 0.75] for the DDB, NL, NL-hyp, SR, CompOSE, CSE, and PWP models, respectively. Beyond 600 MeV·fm-3, the overall correlation declines, reaching 0.8 at 800 MeV·fm-3, while the individual correlations adjust to [0.99, 0.94, 0.87, 0.78, 0.95, 0.37, 0.66] for the same models. As might be expected, the agnostic descriptions show the smallest correlations, with PWP models still showing a correlation above 70%, 75% at MeV·fm-3. At MeV·fm-3, the correlation suffers an overall decrease which is particularly strong for the agnostic descriptions: for CSE it goes below 50%. Given these findings, we applied the symbolic regression method, as described in Sec. 2.2, to further investigate potential relationships between NS maximum mass with pressure and baryon density at a given energy density. This analysis helped us establish a relationship for mixed EoS models at an energy density of 800 MeV fm-3. The obtained equation is
@@
where represents the pressure and is the baryon density at an energy density of 800 MeV.fm-3. In this analysis, we considered a total of 700 EoSs, where 100 EoSs were randomly selected from each of the seven different EoS sets. At a fixed energy density of 800 MeV/fm3, we examined the corresponding baryon density values for all these EoSs. We found that the baryon density spans a range from 0.63 fm-3 (minimum) to 0.77 fm-3 (maximum), indicating noticeable variation across the different models in this high-density regime. This correlation is shown in the right panel of Figure 3. Overall, it illustrates that over 90% of the models within the mixed set adhere closely to the derived equation, highlighting its robustness and practical applicability. The lower panel displays the fit error () as a function of , which remains below 4%, indicating the reliability of our fitted function. This relationship, being largely model-independent and reliable, will be employed for realistic applications in the following.
@@
We used the PyMultiNest sampler [89, 90] with 1500 live points on the Deucalion high performance computing (HPC) system at the University of Minho, Portugal, using 4 nodes with 128 CPUs each. The summary of CPU times for both analyses is shown in Table 1. It shows a dramatic reduction in sampling time for the scenario where the TOV equation was bypassed. This resulted in a reduction in CPU time of more than a factor of seven. We then examined the posterior distributions for the NS mass-radius relationships for both cases. The comparison involved around 8,000 EoSs for each case, as can be seen from Table 1.
@@ 3.1.2 Mass-Radius-Tidal deformability
The equations of state can be immediately obtained, from the observed mass-radius-tidal deformability, by using separately Eqs. (4) & (5)(SRM-MR) and Eqs. (6) & (7)(SRM-M). To test the reliability of these fits, we selected from each of the different sets shown in Figure 2, 100 MR and 100 M curves to reconstruct the EoSs using the relationships in Eqs. (4-7). These 700 reconstructed EoSs are compared to the original EoSs (All) in Figure 5. The 90% confidence interval for the variation of pressure with energy density reconstructed using SRM-MR (light pink) and SRM-M (cyan) are compared to the original EoSs (black). The reconstructed bands show a good overlap with the original one in the range MeV/fm3. It may be noted that the deviation for the reconstructed EoSs increases somewhat with density. We have also repeated the calculations for different sets of randomly selected EoSs, the results remained more-or-less the same.
@@ 4 Conclusions
We have obtained the expressions relating the equations of state to the NS observables such as the radius, and tidal deformability in the mass range of 1-2 . The EoSs derived using these expressions (Eqs. (4-7)) using NS observables are found to have good overlap with the actual EoSs.